OpenStudy (anonymous):

Determine this equation using the laws of exponents: 16x2^x+2

OpenStudy (anonymous):

Is this the correct answer 2^4x+6?

OpenStudy (anonymous):

2^4x2^x+2 2^4x+6

OpenStudy (anonymous):

could you enter that using the equation button below.

OpenStudy (anonymous):

\[16x2^(x+2)\]

OpenStudy (anonymous):



OpenStudy (anonymous):

hmm much better.

OpenStudy (anonymous):

so, \[\Large{16\times2^{x+2}=2^4\times2^{x+2}=2^{4+x+2}=\boxed{2^{x+6}}}\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw.